# Conformal Defects in Neural Network Field Theories

### v1

@@ Abstract
I Introduction
Conformal Field Theories (CFTs) occupy a distinguished place within the space of Quantum Field Theories (QFTs). They are characterized by an enhanced set of (global) spacetime symmetries, furnished by the extension of the Poincaré group to include scaling and special conformal transformations. They are concretely defined in that they lack an intrinsic energy scale, and thus require no such information in their specification. CFTs are often found at fixed points of Renormalisation Group (RG) flows, and are therefore important elements in the study of the behavior of QFTs with relevant couplings. CFTs also exist in arbitrary dimensions, and — in certain, often supersymmetric, cases — can be related across dimensions via compactifications and/or RG flows. Furthermore, unitarity is not a strictly necessary condition for defining a CFT; non-unitary CFTs find application in modeling numerous physical systems and have rich spaces of conformal boundary/defect conditions, see e.g. Chalabi:2022qit ; Herzog:2024zxm for a recent discussion. Because of their ubiquity, CFTs play a crucial role in the study of many areas of physics, from critical phenomena in condensed matter systems to string/M-theory and mathematics; hence, new methods and insights in the understanding of CFTs can often impact the study of disparate fields in surprising ways.
@@ I Introduction
This example of a free field theory in the presence of a boundary can be understood in the NN context through a reduced symmetry in the NN parameters, collectively referred to by $\Theta$ throughout the body of the text, whose distribution $P(\Theta)$. Consider a conformally invariant $d$-dimensional measure $\mathrm{d}P(\Theta)$ in parameter space. Breaking this symmetry down to a (d1)(d-1)-dimensional conformal subgroup can be achieved, for instance, by breaking the translational symmetry along one direction in $\Theta$-space, e.g. by freezing the value of one of the parameters, $\Theta_{\perp}=c$. Indeed, this breaks the symmetry in the $\Theta_{\perp}$ direction while preserving it along the remaining $\Theta\setminus\{\Theta_{\perp}\}$. This logic can be extended to the breaking of other kinds of symmetries, e.g. rotations, and to higher co-dimension defects, in a straightforward fashion.
There is an alternative and particularly intriguing way to think of this symmetry breaking effect. In the NN-FT picture, the input space for the NN is regarded as the background space(-time) on which the field theory is defined; here, for simplicity, we will take this to be $\mathbb{R}^{d}$. The introduction co-dimension $q$ defect from the input space perspective has conceptual overlap with the so-called ‘manifold hypothesis’, which states that data is typically distributed on a submanifold immersed [^1] Note that it is not guaranteed that the intrinsic topology of the ‘data submanifold’ coincides with the (pullback) subset topology in the ambient space. Hence, it is not always the case that there is a smooth embedding. in a higher-dimensional input space; see fefferman2016testing for an introduction to the topic. Thus, one could ask how out-of-distribution or off-of-defect (OOD) samples, i.e. those lying away from the support of the true data distribution, can be used to probe the data submanifold itself. That is, how do expectations of OOD samples with respect to the joint distribution formed by the parameters spanning the directions along the submanifold and the directions normal to it characterize the submanifold structure? Addressing these questions requires methods familiar from the study of correlation functions in defect CFTs.
@@
In the present work, we propose a methodology to construct conformal symmetry breaking layers in NN-FTs and thereby incorporate conformal defects into the general paradigm. We briefly review the necessary components of building conformal fields from NNs and how defects fit into the embedding space formalism in § II. In § III, we synthesize these concepts into a framework for building conformal defects of arbitrary codimension from NNs. With this formalism in hand, we explore two classes of toy models to elucidate its utility in § IV. We then recapitulate our findings and discuss future directions in § V. In App. A, we collect intermediate results that are useful for computing two-point functions in the toy models that we consider.
II Background
@@ II.1 NN-FTs for conformal fields
As a brief aside, we note that global symmetries act differently in the context of NN-FTs since they are transformations in field (or network) space rather than in the input or parameter spaces. For instance, if we consider the analog of an $O(N)$ model of NN-FTs, we would have a collection of architectures $\Phi_{a}(X)$ that are rotated into one another, $\Phi_{b}(X)=R_{b}{}^{a}\Phi_{a}(X)$ by $R\in O(N)$. The concept of global symmetries occurs naturally in the context of defects in NN-FTs, but its appearance is incidental to breakings of the symmetries of the input space which preserve a global subgroup of normal bundle symmetries; we will momentarily forgo further discussion on this topic, only to return to it in § V.
@@
The constant $C_{abc}$ appearing in the three-point function is the operator product expansion (OPE) coefficient. Higher-point functions are less constrained, in that they can determined by kinematics only up to functions of the conformal cross ratios
@@
$\left<\Phi_{a}(x_{1})\Phi_{b}(x_{2})\Phi_{c}(x_{3})\Phi_{d}(x_{4})\right>=\left(\frac{|x_{2}-x_{4}|}{|x_{1}-x_{4}|}\right)^{\Delta_{a}-\Delta_{b}}\left(\frac{|x_{1}-x_{4}|}{|x_{1}-x_{3}|}\right)^{\Delta_{c}-\Delta_{d}}\frac{g(u,v)}{|x_{1}-x_{2}|^{\Delta_{a}+\Delta_{b}}|x_{3}-x_{4}|^{\Delta_{c}+\Delta_{d}}}~.$ (7)
@@
II.2 Conformal defects in embedding space
@@ III Conformal defects in NN-FTs
Consider a conformal field $\Phi(X)$ constructed from an NN acting on an input space spanned by the restriction of the embedding space coordinates $\{X\}$ to a physical point on the Poincaré section. As reviewed in § II.1, one of the constraints imposed on the choice of architecture by the presence of conformal symmetry [^4] While it is not a question that we directly address in the current work, should we allow for the architecture $\Phi_{s_{1}\ldots s_{J}}(X)$ to transform in a non-trivial, symmetric traceless, spin-$J$ representation of $SO(d)\subset SO(d+1,1)$, i.e. a spinning conformal field, there are additional constraints afforded by the conformal symmetry group in the form of transversality with respect to a set of auxiliary vectors $z^{s_{i}}$ encoding the spin of the field, e.g. $z^{s_{1}}\Phi_{s_{1}\ldots s_{J}}=0$. We will take up this question in future work. is that $\Phi(X)\equiv\Phi_{\Delta}(X)$ must be a homogeneous function of $X$ with scaling weight $\Delta$, i.e. $\Phi_{\Delta}(X)\to\Phi_{\Delta}(\lambda X)=\lambda^{-\Delta}\Phi_{\Delta}(X)$ under scale transformations $X\to\lambda X$. Since our aim is to preserve a $p$-dimensional conformal subgroup, with $p<d$, we require that any conformal defect constructed from an NN must similarly be a homogeneous function of the subset of inputs spanning $\mathfrak{M}_{p}\hookrightarrow\mathbb{R}^{d}$ Billo:2016cpy . Further, if the additional normal bundle symmetry is preserved, then we also require that the network architecture transforms in an irreducible representation labeled by the ‘transverse spin’ $s$.
@@ IV.1 Monomial NN-FTs
$\langle\Phi_{\alpha}(X)\rangle$ $=\mathbb{E}_{P(\Theta)}\left[\Phi_{\alpha}(X)\right]=\sum_{\widehat{\alpha}=0}^{\alpha}\frac{\Gamma(\alpha+1)}{\Gamma(\widehat{\alpha}+1)\Gamma(\alpha-\widehat{\alpha}+1)}\mathbb{E}_{\widehat{P}(\Theta)}\left[(X\bullet\Theta)^{\widehat{\alpha}}\right]\mathbb{E}_{\widetilde{P}(\Theta)}\left[(X\circ\Theta)^{\alpha-\widehat{\alpha}}\right$]
@@
In moving from the first to the second line, we note that the expectation values $\mathbb{E}[\ldots]$ are only non-vanishing if $\widehat{\alpha}\in 2\mathbb{N}$ and $\alpha-\widehat{\alpha}\in 2\mathbb{N}$. In the final equality, we restrict to the Poincaré section. In contrast to the defect-localized field $\widehat{\varphi}(X)$, the ambient field $\Phi_{\alpha}(X)$ is not restricted to the submanifold $X\circ X\overset{\rm P.S.}{=}0$, and we see that the one-point function of this defect architecture is precisely of the form that one would expect for the one-point function of an ambient scalar of dimension $\alpha=-\Delta$ in the presence of a conformal defect Billo:2016cpy . One further point to note is that the overall normalization vanishes when $\widehat{\sigma}=\widetilde{\sigma}$: in that case, the defect becomes trivial.
In the binomial expansion in the first line of eq. (IV.1), we see a structure similar to that of the defect OPE. Indeed, looking back to eq. (13) and dropping the towers of descendants ($n=0$), we can draw a direct analogy with
@@
$\mathbb{E}\left[\Phi_{\alpha_{1}}(X_{1})\widehat{\varphi}_{\widehat{\alpha}_{2}}(X_{2})\right]=\sum_{\widehat{\alpha}_{1}=0}^{\alpha_{1}}\frac{\Gamma(\alpha_{1}+1)}{\Gamma(\widehat{\alpha}_{1}+1)\Gamma(\alpha_{1}-\widehat{\alpha}_{1}+1)}\mathbb{E}_{\widehat{P}(\Theta)}\left[(X_{1}\bullet\Theta)^{\widehat{\alpha}_{1}}(X_{2}\bullet\Theta)^{\widehat{\alpha}_{2}}\right]\mathbb{E}_{\widetilde{P}(\Theta)}\left[(X_{1}\circ\Theta)^{\alpha_{1}-\widehat{\alpha}_{1}}\right]~.$ (25)
@@
$\mathbb{E}[\Phi_{\alpha_{1}}(X_{1})\Phi_{\alpha_{2}}(X_{2})]=$ $\frac{\Gamma(\alpha_{2}+1)}{2^{\frac{\alpha_{2}+\alpha_{1}}{2}}\Gamma(\frac{\alpha_{1}-\alpha_{2}+2}{2})}(\widetilde{\sigma}^{2}-\widehat{\sigma}^{2})^{\frac{\alpha_{1}-\alpha_{2}}{2}}\|X_{1}\|_{\circ}^{\alpha_{1}}\|X_{2}\|_{\circ}^{\alpha_{2}}\left(2\widetilde{\sigma}^{2}\cos\psi-\widehat{\sigma}^{2}\chi\right)^{\alpha_{2}}\times$ (32)
$\qquad\times{}_{2}F_{1}\left(\frac{1-\alpha_{2}}{2},-\frac{\alpha_{2}}{2},\frac{\alpha_{1}-\alpha_{2}+2}{2},\frac{4(\widetilde{\sigma}^{2}-\widehat{\sigma}^{2})^{2}}{(2\widetilde{\sigma}^{2}\cos\psi-\widehat{\sigma}^{2}\chi)^{2}}\right),$
@@ IV.2 Reciprocal NN-FTs
we can shift the divergence to the boundary of the $u$-integration region. In principle, further care must be taken upon treating eq. (46), as it is only valid when all monomials $\Theta\bullet X_{k}$ have the same sign. Nevertheless, as we will discuss shortly, we can adopt a regularisation scheme which, by suitably removing all divergences, will allow us to forgo this condition.
@@
$G^{(n)}(\Delta_{1},\ldots,\Delta_{n}):=\frac{\Gamma(\Delta_{1}+\cdots+\Delta_{n})}{\Gamma(\Delta_{1})\cdots\Gamma(\Delta_{n})}\sqint\,\mathrm{d}^{n}u\int\mathrm{d}^{p+2}\Theta~\widehat{P}(\Theta)\frac{\delta(1-\sum_{k=1}^{n}u_{k})\prod_{k=1}^{n}u^{\Delta_{k}-1}_{k}}{(\sum_{k=1}^{n}u_{k}(X_{k}\bullet\Theta))^{\sum_{k=1}^{n}\Delta_{k}}}~,$ (47)
@@
$G^{(n)}(\Delta_{1},\ldots,\Delta_{n})$ $=\frac{1}{\sqrt{\pi}}\Gamma\left(\frac{1-\Delta}{2}\right)\frac{\Gamma(\Delta_{1}+\cdots+\Delta_{n})}{\Gamma(\Delta_{1})\cdots\Gamma(\Delta_{n})}\times$
$\quad\times\sqint\,\mathrm{d}^{n}u~\delta\left(1-\sum_{k=1}^{n}u_{k}\right)\prod_{k=1}^{n}u^{\Delta_{k}-1}_{k}(2\hat{\sigma}^{2}Z(u)\bullet Z(u))^{-\frac{\Delta}{2}},$ (50)
@@
Let us illustrate this cut-off regularisation explicitly in the case of two $u$-integrals, corresponding to correlators between two defect insertions. For now, let $I(\Delta_{1},\Delta_{2})$ denote the $u$-integral part of $G^{(2)}(\Delta_{1},\Delta_{2})$,
@@
$=\sqint\mathrm{d}u~u^{\Delta_{1}-1}(1-u)^{\Delta_{2}-1}\left(u^{2}X_{1}\bullet X_{1}+2u(1-u)X_{1}\bullet X_{2}+(1-u)^{2}X_{2}\bullet X_{2}\right)^{-\frac{\Delta_{1}+\Delta_{2}}{2}}.$ (53)
@@
Having properly defined the integration of defect fields $\widehat{\varphi}$ in eq. (IV.2), we can evaluate the various correlators of ambient and defect fields, $\Phi$ and $\widehat{\varphi}$. The one-point correlation functions of defect fields $\widehat{\varphi}$ directly follow from the definition of $G^{(1)}$. Furthermore, one finds that for all $\Delta>0$, these vanish, as expected for any defect conformal primary (c.f. eq. (11)). The expansion of the ambient field in terms of normal and defect ones in eq. (44) informs us that the ambient field one-point function also vanishes,
@@
This is may appear to be an amusing configuration, in that the coupling to the defect identity vanishes, but it reflects the observation that the expansion never involves the identity defect field.
@@
Consider now the three classes of two-point functions $\mathbb{E}[\widehat{\varphi}_{\widehat{\Delta}_{1}}(X_{1})\widehat{\varphi}_{\widehat{\Delta}_{2}}(X_{2})]$, $\mathbb{E}[\Phi_{\Delta}(X_{1})\widehat{\varphi}_{\widehat{\Delta}}(X_{2})]$, $\mathbb{E}[\Phi_{\Delta_{1}}(X_{1})\Phi_{\Delta_{2}}(X_{2})]$. The two-point function between defect fields $\widehat{\varphi}_{\widehat{\Delta}}$ is, by definition, the integral $G^{(2)}$ defined in eq. (IV.2). A detailed calculation of the $u$-integral was given in the previous subsection. The full correlator for $\Delta\in\mathbb{R}_{+}$ reads
@@
Before addressing the issue concerning reflection positivity, let us evaluate the mixed ambient-defect two-point function, which requires two separate steps. First, one expands the ambient field as in eq. (44) in terms of defect fields $\widehat{\varphi}$. The mixed correlator becomes an infinite sum of ambient one-point and defect two-point functions. For the former we use eq. (IV.1) (or equivalently eq. (49)), while the latter is given in eq. (57). The resulting sum collapses down to
$\mathbb{E}[\Phi_{\Delta}(X_{1})\widehat{\varphi}_{\widehat{\Delta}_{2}}(X_{2})]$ $=\frac{2^{\frac{\Delta_{1}-{\widehat{\Delta}_{2}}}{2}}\sec(\pi{\widehat{\Delta}_{2}})}{\Gamma(\Delta_{1})\Gamma(\frac{{\widehat{\Delta}_{2}}-\Delta_{1}}{2}+1)}\tilde{\sigma}^{{\widehat{\Delta}_{2}}-\Delta_{1}}\|X_{1}\|_{\circ}^{{\widehat{\Delta}_{2}}-\Delta_{1}}(\hat{\sigma}^{2}X_{1}\bullet X_{2})^{-{\widehat{\Delta}_{2}}}~,\qquad{\widehat{\Delta}_{2}}\geq\Delta_{1}~,$ (58)
and vanishes when $\Delta_{1}>{\widehat{\Delta}_{2}}$. Again, this has the structural form expected for a mixed ambient-defect correlator respecting defect conformal symmetry. Also, we note that taking the ambient field to be the identity, and therefore $\Delta_{1}\to 0$, recovers the vanishing of the one-point function for defect fields, and trivially also recovers the vanishing of the one-point function for ambient fields for any $\Delta_{1}>0$ when taking ${\widehat{\Delta}_{2}}\to 0$.

### v2

# Southampton, SO17 1BJ, UK

@@ Abstract
arxiv: 2512.07946
1 Introduction
Conformal Field Theories (CFTs) occupy a distinguished place within the space of Quantum Field Theories (QFTs). They are characterized by an enhanced set of (global) spacetime symmetries, furnished by the extension of the Poincaré group to include scaling and special conformal transformations. They are concretely defined in that they lack an intrinsic energy scale, and thus require no such information in their specification. CFTs are often found at fixed points of Renormalization Group (RG) flows, and are therefore important elements in the study of the behavior of QFTs with relevant couplings. CFTs also exist in arbitrary dimensions, and — in certain, often supersymmetric, cases — can be related across dimensions via compactifications and/or RG flows. Furthermore, unitarity is not a strictly necessary condition for defining a CFT; non-unitary CFTs find application in modeling numerous physical systems and have rich spaces of conformal boundary/defect conditions, see e.g. Chalabi:2022qit ; Herzog:2024zxm for a recent discussion. Because of their ubiquity, CFTs play a crucial role in the study of many areas of physics, from critical phenomena in condensed matter systems to string/M-theory and mathematics; hence, new methods and insights in the understanding of CFTs can often impact the study of disparate fields in surprising ways.
@@ 1 Introduction
This example of a free field theory in the presence of a boundary can be understood in the NN context through a reduced symmetry in the NN parameters, collectively referred to by $\Theta$ throughout the body of the text, whose distribution is $P(\Theta)$. Consider a conformally invariant $d$-dimensional measure $\mathrm{d}P(\Theta)$ in parameter space. Breaking this symmetry down to a (d1)(d-1)-dimensional conformal subgroup can be achieved, for instance, by breaking the translational symmetry along one direction in $\Theta$-space, e.g. by freezing the value of one of the parameters, $\Theta_{\perp}=c$. Indeed, this breaks the symmetry in the $\Theta_{\perp}$ direction while preserving it along the remaining $\Theta\setminus\{\Theta_{\perp}\}$. This logic can be extended to the breaking of other kinds of symmetries, e.g. rotations, and to higher co-dimension defects, in a straightforward fashion.
There is an alternative and particularly intriguing way to think of this symmetry breaking effect. In the NN-FT picture, the input space for the NN is regarded as the background space(-time) on which the field theory is defined; here, for simplicity, we will take this to be $\mathbb{R}^{d}$. The introduction of a co-dimension $q$ defect from the input space perspective has conceptual overlap with the so-called ‘manifold hypothesis’, which states that data is typically distributed on a submanifold immersed [^1] Note that it is not guaranteed that the intrinsic topology of the ‘data submanifold’ coincides with the (pullback) subset topology in the ambient space. Hence, it is not always the case that there is a smooth embedding. in a higher-dimensional input space; see fefferman2016testing for an introduction to the topic. Thus, one could ask how out-of-distribution or off-of-defect (OOD) samples, i.e. those lying away from the support of the true data distribution, can be used to probe the data submanifold itself. That is, how do expectations of OOD samples with respect to the joint distribution formed by the parameters spanning the directions along the submanifold and the directions normal to it characterize the submanifold structure? Addressing these questions requires methods familiar from the study of correlation functions in defect CFTs.
@@
In the present work, we propose a methodology to construct conformal symmetry breaking layers in NN-FTs and thereby incorporate conformal defects into the general paradigm. We briefly review the necessary components of building conformal fields from NNs and how defects fit into the embedding space formalism in § 2. In § 3, we synthesize these concepts into a framework for building conformal defects of arbitrary codimension from NNs. With this formalism in hand, we explore two classes of toy models to elucidate its utility in § 4. We then recapitulate our findings and discuss future directions in § 5. In App. A, we collect intermediate results that are useful for computing two-point functions in the toy models that we consider.
2 Background
@@ 2.1 NN-FTs for conformal fields
As a brief aside, we note that global symmetries act differently in the context of NN-FTs since they are transformations in field (or network) space rather than in the input or parameter spaces. For instance, if we consider the analog of an $O(N)$ model of NN-FTs, we would have a collection of architectures $\Phi_{a}(X)$ that are rotated into one another, $\Phi_{b}(X)=R_{b}{}^{a}\Phi_{a}(X)$ by $R\in O(N)$. The concept of global symmetries occurs naturally in the context of defects in NN-FTs, but its appearance is incidental to breakings of the symmetries of the input space which preserve a global subgroup of normal bundle symmetries; we will momentarily forgo further discussion on this topic, only to return to it in § 5.
@@
The constant $C_{abc}$ appearing in the three-point function is the operator product expansion (OPE) coefficient. Higher-point functions are less constrained, in that they can be determined by kinematics only up to functions of the conformal cross ratios
@@
$\left<\Phi_{a}(x_{1})\Phi_{b}(x_{2})\Phi_{c}(x_{3})\Phi_{d}(x_{4})\right>=\frac{\left(\frac{|x_{2}-x_{4}|}{|x_{1}-x_{4}|}\right)^{\Delta_{a}-\Delta_{b}}\left(\frac{|x_{1}-x_{4}|}{|x_{1}-x_{3}|}\right)^{\Delta_{c}-\Delta_{d}}}{|x_{1}-x_{2}|^{\Delta_{a}+\Delta_{b}}|x_{3}-x_{4}|^{\Delta_{c}+\Delta_{d}}}g(u,v)~.$ (7)
@@
2.2 Conformal defects in embedding space
@@ 3 Conformal defects in NN-FTs
Consider a conformal field $\Phi(X)$ constructed from an NN acting on an input space spanned by the restriction of the embedding space coordinates $\{X\}$ to a physical point on the Poincaré section. As reviewed in § 2.1, one of the constraints imposed on the choice of architecture by the presence of conformal symmetry [^4] While it is not a question that we directly address in the current work, should we allow for the architecture $\Phi_{s_{1}\ldots s_{J}}(X)$ to transform in a non-trivial, symmetric traceless, spin-$J$ representation of $SO(d)\subset SO(d+1,1)$, i.e. a spinning conformal field, there are additional constraints afforded by the conformal symmetry group in the form of transversality with respect to a set of auxiliary vectors $z^{s_{i}}$ encoding the spin of the field, e.g. $z^{s_{1}}\Phi_{s_{1}\ldots s_{J}}=0$. We will take up this question in future work. is that $\Phi(X)\equiv\Phi_{\Delta}(X)$ must be a homogeneous function of $X$ with scaling weight $\Delta$, i.e. $\Phi_{\Delta}(X)\to\Phi_{\Delta}(\lambda X)=\lambda^{-\Delta}\Phi_{\Delta}(X)$ under scale transformations $X\to\lambda X$. Since our aim is to preserve a $p$-dimensional conformal subgroup, with $p<d$, we require that any conformal defect constructed from an NN must similarly be a homogeneous function of the subset of inputs spanning $\mathfrak{M}_{p}\hookrightarrow\mathbb{R}^{d}$ Billo:2016cpy . Further, if the additional normal bundle symmetry is preserved, then we also require that the network architecture transforms in an irreducible representation labeled by the ‘transverse spin’ $s$.
@@ 4.1 Monomial NN-FTs
$\langle\Phi_{\alpha}(X)\rangle$ $=\mathbb{E}_{P(\Theta)}\left[\Phi_{\alpha}(X)\right]$
$=\sum_{\widehat{\alpha}=0}^{\alpha}\frac{\Gamma(\alpha+1)}{\Gamma(\widehat{\alpha}+1)\Gamma(\alpha-\widehat{\alpha}+1)}\mathbb{E}_{\widehat{P}(\Theta)}\left[(X\bullet\Theta)^{\widehat{\alpha}}\right]\mathbb{E}_{\widetilde{P}(\Theta)}\left[(X\circ\Theta)^{\alpha-\widehat{\alpha}}\right$]
@@
In moving from the second to the third line, we note that the expectation values $\mathbb{E}[\ldots]$ are only non-vanishing if $\widehat{\alpha}\in 2\mathbb{N}$ and $\alpha-\widehat{\alpha}\in 2\mathbb{N}$. In the final equality, we restrict to the Poincaré section. In contrast to the defect-localized field $\widehat{\varphi}(X)$, the ambient field $\Phi_{\alpha}(X)$ is not restricted to the submanifold $X\circ X\overset{\rm P.S.}{=}0$, and we see that the one-point function of this defect architecture is precisely of the form that one would expect for the one-point function of an ambient scalar of dimension $\alpha=-\Delta$ in the presence of a conformal defect Billo:2016cpy . One further point to note is that the overall normalization vanishes when $\widehat{\sigma}=\widetilde{\sigma}$: in that case, the defect becomes trivial.
In the binomial expansion in the first line of eq. (4.1), we see a structure similar to that of the defect OPE. Indeed, looking back to eq. (13) and dropping the towers of descendants ($n=0$), we can draw a direct analogy with
@@
$\hskip-28.45274pt\mathbb{E}\left[\Phi_{\alpha_{1}}(X_{1})\widehat{\varphi}_{\widehat{\alpha}_{2}}(X_{2})\right]$ $=\sum_{\widehat{\alpha}_{1}=0}^{\alpha_{1}}\frac{\Gamma(\alpha_{1}+1)}{\Gamma(\widehat{\alpha}_{1}+1)\Gamma(\alpha_{1}-\widehat{\alpha}_{1}+1)}\mathbb{E}_{\widehat{P}(\Theta)}\left[(X_{1}\bullet\Theta)^{\widehat{\alpha}_{1}}(X_{2}\bullet\Theta)^{\widehat{\alpha}_{2}}\right]\times$
$\qquad\qquad\times\mathbb{E}_{\widetilde{P}(\Theta)}\left[(X_{1}\circ\Theta)^{\alpha_{1}-\widehat{\alpha}_{1}}\right]~.$ (25)
@@
$\mathbb{E}[\Phi_{\alpha_{1}}(X_{1})\Phi_{\alpha_{2}}(X_{2})]=\frac{\Gamma(\alpha_{2}+1)}{2^{\frac{\alpha_{2}+\alpha_{1}}{2}}\Gamma(\frac{\alpha_{1}-\alpha_{2}+2}{2})}(\widetilde{\sigma}^{2}-\widehat{\sigma}^{2})^{\frac{\alpha_{1}-\alpha_{2}}{2}}\|X_{1}\|_{\circ}^{\alpha_{1}}\|X_{2}\|_{\circ}^{\alpha_{2}}\times$ (32)
$\times\left(2\widetilde{\sigma}^{2}\cos\psi-\widehat{\sigma}^{2}\chi\right)^{\alpha_{2}}{}_{2}F_{1}\left(\frac{1-\alpha_{2}}{2},-\frac{\alpha_{2}}{2},\frac{\alpha_{1}-\alpha_{2}+2}{2},\frac{4(\widetilde{\sigma}^{2}-\widehat{\sigma}^{2})^{2}}{(2\widetilde{\sigma}^{2}\cos\psi-\widehat{\sigma}^{2}\chi)^{2}}\right),$
@@ 4.2 Reciprocal NN-FTs
we can shift the divergence to the boundary of the $u$-integration region. In principle, further care must be taken upon treating eq. (46), as it is only valid when all monomials $\Theta\bullet X_{k}$ have the same sign. Nevertheless, as we will discuss shortly, we can adopt a regularization scheme which, by suitably removing all divergences, will allow us to forgo this condition.
@@
$G^{(n)}(\Delta_{1},\ldots,\Delta_{n}):=\frac{\Gamma(\sum_{i=1}^{n}\Delta_{i})}{\prod_{i=1}^{n}\Gamma(\Delta_{i})}\sqint\,\mathrm{d}^{n}u\int\mathrm{d}^{p+2}\Theta~\widehat{P}(\Theta)\frac{\delta(1-\sum_{k=1}^{n}u_{k})\prod_{k=1}^{n}u^{\Delta_{k}-1}_{k}}{(\sum_{k=1}^{n}u_{k}(X_{k}\bullet\Theta))^{\sum_{k=1}^{n}\Delta_{k}}}~,$ (47)
@@
$G^{(n)}(\Delta_{1},\ldots,\Delta_{n})$ $=\frac{1}{\sqrt{\pi}}\Gamma\left(\frac{1-\Delta}{2}\right)\frac{\Gamma(\Delta_{1}+\cdots+\Delta_{n})}{\Gamma(\Delta_{1})\cdots\Gamma(\Delta_{n})}\times$ (50)
$\quad\times\sqint\,\mathrm{d}^{n}u~\delta\left(1-\sum_{k=1}^{n}u_{k}\right)\prod_{k=1}^{n}u^{\Delta_{k}-1}_{k}(2\hat{\sigma}^{2}Z(u)\bullet Z(u))^{-\frac{\Delta}{2}},$
@@
Let us illustrate this cut-off regularization explicitly in the case of two $u$-integrals, corresponding to correlators between two defect insertions. For now, let $I(\Delta_{1},\Delta_{2})$ denote the $u$-integral part of $G^{(2)}(\Delta_{1},\Delta_{2})$,
@@
$=\sqint\mathrm{d}u~u^{\Delta_{1}-1}(1-u)^{\Delta_{2}-1}\times$ (53)
$\quad\times\left(u^{2}X_{1}\bullet X_{1}+2u(1-u)X_{1}\bullet X_{2}+(1-u)^{2}X_{2}\bullet X_{2}\right)^{-\frac{\Delta_{1}+\Delta_{2}}{2}}.$
@@
Having properly defined the integration of defect fields $\widehat{\varphi}$ in eq. (50), we can evaluate the various correlators of ambient and defect fields, $\Phi$ and $\widehat{\varphi}$. The one-point correlation functions of defect fields $\widehat{\varphi}$ directly follow from the definition of $G^{(1)}$. Furthermore, one finds that for all $\Delta>0$, these vanish, as expected for any defect conformal primary (c.f. eq. (11)). The expansion of the ambient field in terms of normal and defect ones in eq. (44) informs us that the ambient field one-point function also vanishes,
@@
This may appear to be an amusing configuration, in that the coupling to the defect identity vanishes, but it reflects the observation that the expansion never involves the identity defect field.
@@
Consider now the three classes of two-point correlation functions: $\mathbb{E}[\widehat{\varphi}_{\widehat{\Delta}_{1}}(X_{1})\widehat{\varphi}_{\widehat{\Delta}_{2}}(X_{2})]$, $\mathbb{E}[\Phi_{\Delta}(X_{1})\widehat{\varphi}_{\widehat{\Delta}}(X_{2})]$, and $\mathbb{E}[\Phi_{\Delta_{1}}(X_{1})\Phi_{\Delta_{2}}(X_{2})]$. The two-point function between defect fields $\widehat{\varphi}_{\widehat{\Delta}}$ is, by definition, the integral $G^{(2)}$ defined in eq. (50). A detailed calculation of the $u$-integral was given in the previous subsection. The full correlator for $\Delta\in\mathbb{R}_{+}$ reads
@@
Before addressing the issue concerning reflection positivity, let us evaluate the mixed ambient-defect two-point function, which requires two separate steps. First, one expands the ambient field as in eq. (44) in terms of defect fields $\widehat{\varphi}$. The mixed correlator becomes an infinite sum of ambient one-point and defect two-point functions. For the former we use eq. (4.1) (or equivalently eq. (49)), while the latter is given in eq. (57). The resulting sum collapses down to
$\mathbb{E}[\Phi_{\Delta}(X_{1})\widehat{\varphi}_{\widehat{\Delta}_{2}}(X_{2})]$ $=\frac{2^{\frac{\Delta_{1}-{\widehat{\Delta}_{2}}}{2}}\sec(\pi{\widehat{\Delta}_{2}})}{\Gamma(\Delta_{1})\Gamma(\frac{{\widehat{\Delta}_{2}}-\Delta_{1}}{2}+1)}\tilde{\sigma}^{{\widehat{\Delta}_{2}}-\Delta_{1}}\|X_{1}\|_{\circ}^{{\widehat{\Delta}_{2}}-\Delta_{1}}(\hat{\sigma}^{2}X_{1}\bullet X_{2})^{-{\widehat{\Delta}_{2}}}~,$ (58)
when ${\widehat{\Delta}_{2}}\geq\Delta_{1}$ and vanishes otherwise. Again, this has the structural form expected for a mixed ambient-defect correlator respecting defect conformal symmetry. Also, we note that taking the ambient field to be the identity, and therefore $\Delta_{1}\to 0$, recovers the vanishing of the one-point function for defect fields, and trivially also recovers the vanishing of the one-point function for ambient fields for any $\Delta_{1}>0$ when taking ${\widehat{\Delta}_{2}}\to 0$.